Circular No. 8680

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2006R–2006T, 2006V, 2006X, 2006al, 2006am, 2006an

Two new supernovae have been reported from unfiltered CCD images: 2006am in NGC 5630 by E. Lee and W. Li (LOSS/KAIT; cf. IAUC 8669), and 2006an by R. Quimby and F. Castro, University of Texas (0.45-m ROTSE-IIIb telescope; cf. IAUC 8622).

SN  2006 UT  α_{2000}  δ_{2000}  Mag.  Offset
2006am Feb. 22.53 14^{h}27^{m}37.^{s}24 +41^{o}15′35′′.4 18.5 7′′.5 E, 6′′.8 N
2006an Feb. 21.27 12 14 38.68 +12 13 46.9 17.9 1′′ W, 1′′ S

Additional KAIT magnitudes for 2006am: 2005 July 4.21 UT, [19.5; Feb. 23.55, 18.5; 24.54, 18.5. SN 2005dp also appeared in NGC 5630. A coaddition of images taken between 2004 Dec. 15 and 2005 June 8 shows nothing at the location of 2006an, with limiting mag ∼ 19.0; 2006an was also detected on 2006 Feb. 26.27 at mag ∼ 17.7. The positional uncertainty for 2006an above is estimated as ± 0′′.6; the specified offset is with respect to a galaxy that is not significantly detected in the ROTSE-IIIb data, but it appears at mag r = 21.9 in SDSS data with position end figures 38^{s}.73, 47′′.7. SN 2006an is a type-Ia supernova (cf. CBET 413).

S. Blondin, M. Modjaz, R. Kirshner, and P. Challis, Harvard-Smith-sonian Center for Astrophysics, report that a spectrogram (range 350–740 nm) of SN 2006am, obtained on Feb. 27.49 UT by P. Berlind with the Mt. Hopkins 1.5-m telescope (+ FAST), shows it to be a type-IIn supernova. The spectrum consists of a blue continuum and Balmer emission lines, with a narrow component of ∼ 300–400 km/s (FWHM) superposed on a weaker and broader component of ∼ 2000 km/s (FWHM), after removing the NED recession velocity of 2694 km/s for the host galaxy (from The

Updated Zwicky Catalog, http://tdc-www.harvard.edu/uzc/). SN 2006al

is a type-Ia supernova (CBET 411). SNe 2006R and 2006S (cf. IAUC 8664) were found to be type-Ia supernovae within a week of maximum on Jan. 28–29 (CBET 382, 383). SN 2006T (cf. IAUC 8666) was found on Feb. 2 to be an apparently young type-II/IIb event (CBET 386); SN 2006V (cf. IAUC 8667) was a type-II supernova past maximum on Feb. 7 (CBET 392).

A spectrogram of SN 2006X (cf. IAUC 8667) taken on Feb. 8.35 UT shows the object to be an early type-Ia supernova (cf. R. Quimby et al., CBET 393). A ROTSE-IIIb telescope image yielded the following position for 2006X: α = 12^{h}22^{m}53^{s}.88, δ = +15^{o}48′31′′.9 (equinox 2000.0; uncertainty ± 0′′.6). Approximate ROTSE-IIIb unfiltered CCD magnitudes for 2006X: 2005 June 8, [19.0; 2006 Feb. 1.332, [17.9; 7.31, 15.3.

2006 February 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green